40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Feb 6 15:11:21 2013, Jenne, Update, Electronics, "Temporary" power supply situation 
    Reply  Wed Feb 6 15:39:19 2013, Steve, Update, Electronics, DC power supplies in cabinets IMG_0066_1.JPG
       Reply  Wed Feb 6 20:00:06 2013, Manasa, Update, Electronics, BNC cables piled up at every corner 109393779.png
Message ID: 8011     Entry time: Wed Feb 6 15:11:21 2013     Reply to this: 8013
Author: Jenne 
Type: Update 
Category: Electronics 
Subject: "Temporary" power supply situation 

[Jenne, Yuta, Rana, Steve, Manasa]

We have taken stock of the lab "temporary" power supply situation, and things look much better.

This morning, I removed 2 unused power supplies and a function generator from the PSL table - these had been used for MC ringdown things.

I also removed the non-connected cables that had been used for the RAMMON setup, and the EOM temperature controller circuit.

This afternoon, Yuta removed the 2 HV power supplies that were used to keep PZT2 working near the end of its life.  Since we now have the active TTs, these have been turned off for a while, and just needed to be removed.

Manasa removed the power supply under the POP/POX table that was powering the amplifier for POP22.  If we are going to continue using a Thorlabs PD for POP22, then we need to make a twisted pair of wires (~20 feet) to get power from 1X1.  If we are going to (finally) install a gold RF PD, then that will not be necessary.

I removed the power supply sitting near the bottom of the LSC rack, for another amplifier for POP22 (with minicircuits filters attached).  Again, if we get a gold RF PD, we can remove the filters and amplifier.  If not, we can use the existing twisted pair of wires, and plug them into the rack rather than a power supply. 

The power supply under the NE corner of the PSL table was no longer in use.  It was most recently used for amplifiers for the green beat PDs, as Yuta mentioned in elog 6862, those were moved over to 1X2.  In elog 8008 I mentioned that Yuta and I moved those amplifiers over to rack power.

The HV supply, the function generator and the OSA controllers that were on top of the short OMC rack next to the AS table have all been removed.  We need to come up with a better place for the OSAs, since we need to re-install them.  The power supply and the function generator (which was used just for a voltage offset) were formerly used for the output steering PZTs, but lately we have just been using those mirrors as fixed mirrors, since we don't need to steer into the OMC.  Some day, we will replace those mirrors with the output steering active tip tilts, and re-commission the OMC....someday.

The power supply for the amplifier set (that goes with the set of minicircuits filters) for the RAMMON PD (which took light from the IPPOS path) has been removed.  If we determine that we need RAMMON back, we will have to make a twisted pair to power those amplifiers.

SUMMARY:

* If we don't install a gold RF PD for POP22, we need a 20ft twisted pair for +15/GND.

* Also, if we don't install a gold RF PD for POP22, we need to plug the amplifier at the LSC rack into the rack power (twisted pair already exists).

* If we need RAMMON back, we will need a twisted pair to power those amplifiers.

* All other power supplies have been removed, and put away.  We currently have 0 "temporary" power supplies in use in the lab!

ELOG V3.1.3-